# Personalizing story generation by inferring author styles

### v1

@@ 5 Experiments
In this section, we describe the implementation details and evaluation methods.
@@ 5.1 Implementation Details
We prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022; Shashidhar et al., 2024) with a temperature of 0.0 and a maximum token limit of 4096 to generate story rules for the Delta and Oracle methods, as well as for generating the Author Writing Sheet, Author Writing Summary, and their respective persona descriptions (LLMpersona) and personalized story rules (LLMrule). For story generation (LLMstory) across all methods, including Average Author, we evaluate three models: GPT-4o (Bubeck et al., 2023), Llama 3.1 8B, and Llama 3.1 70B (Dubey et al., 2024), using a temperature of 0.7 and top_p of 0.95 (Wang et al., 2024) (See Appendix F for results using Llama models for story generation). Additionally, we implement an ablation variant of our proposed method for Writing Sheet and Writing Summary that excludes the persona description from the system prompt called Writing Sheet nP and Writing Summary nP respectively.
5.2 Automatic Evaluation
We evaluate the performance of personalized story generation using two automated evaluation methods: First, we use GPT-4o-as-a-judge which is shown to be a scalable way to approximate human judgments for complex open-ended tasks (Zheng et al., 2023) and calculate the win-rate of our personalization methods against the non-personalized method, Average Author on two aspects:
  • Faithfulness to Writing History: We evaluate how well the generated story aligns with the author history by measuring win-rates using Claims from the Author Writing Summary as the ground-truth reference for each narrative category (Wang et al., 2024, 2023; Yunusov et al., 2024).
@@ 5.2 Automatic Evaluation
For both aspects, we prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022) to assign scores (1–5) to stories. We randomly shuffle the order of compared pairs, one generated by a personalization method and the other by the Average Author method, to avoid biases. Scores are broken down into four narrative categories: Plot, Creativity, Development, and Language Use (Saha et al., 2024). The story with the higher score in a category is declared the winner, and the overall winner is based on the highest total score across all categories. Additionally, we evaluate personalization using traditional metrics (Xie et al., 2023) including lexical overlap, story diversity, and stylistic similarity (see Appendix D.2 for more details).
5.3 Human Evaluation
We conduct a human evaluation on a subset of our data to further assess subtext and identify any interesting insights that the LLM judges might have missed (Chakrabarty et al., 2024; Subbiah et al., 2024b), by comparing stories generated by personalization methods (Delta, Writing Sheet, Writing Summary) against the Average Author method for similarity to the ground-truth author story (Lyu et al., 2024). We ask annotators to justify their choices after making the comparison. Three annotators, recruited via Upwork (\$17/hour), evaluated 45 author stories in total (15 common to establish agreement and 10 exclusive to each annotator to increase coverage of annotation (Song et al., 2024)). Each author story had three pairs, each containing a story from one personalization method, resulting in 135 annotated story pairs.
@@ 6 Results and Discussion
We now discuss our experimental results and findings.
Table 2: Percentage win-rate for Faithfulness to Writing History evaluated by GPT-4o. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder. Best win-rates for each source is underlined.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 52-32 31-54 49-41 74-20 80-10 68-21 72-19
Reddit 74-21 35-56 58-26 86-7 89-7 74-19 72-19
Storium 50-40 42-45 45-42 68-15 75-8 70-20 68-20
N.Magazine 71-14 57-21 50-36 79-7 71-21 71-21 71-0
New Yorker 53-40 27-67 53-40 67-20 80-20 67-27 80-13
Overall 60-29 38-49 51-37 75-14 79-13 70-22 73-14
Table 3: Percentage win-rate for Similarity with Author Story evaluated by GPT-4o and Human (last column). Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder. Best win-rates for each source is underlined.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP Human Eval
AO3 71-18 30-54 45-32 40-41 60-30 46-38 45-40 37-33
Reddit 91-5 37-46 49-35 54-35 49-42 47-40 42-44 52-37
Storium 60-30 45-40 35-35 52-32 50-35 57-14 14-29 30-30
N.Magazine 64-21 43-36 36-43 50-21 57-21 43-43 50-29 48-30
New Yorker 67-13 27-47 40-40 27-60 27-47 40-60 47-40 70-15
Overall 71-17 36-45 41-37 45-38 49-35 47-39 40-36 47-29
6.1 Automatic Evaluation
Our automatic evaluation results show that, first, Writing Sheet and Summary consistently outperform other methods. Second, persona descriptions improve faithfulness to writing history but improve similarity to the ground truth only for select sources like Reddit. Third, personalization is more effective for sources like Reddit than AO3 and Storium, and for narrative categories like Creativity, and Language Use, compared to Plot.
6.1.1 Faithfulness to Writing History
@@
Table 2 shows results for faithfulness to writing history using GPT-4o as LLMstory for our personalization methods. We observe the following:
Writing Sheet and Summary achieve best faithfulness to writing history:
Writing Sheet and Summary achieve the highest win-rates across all sources, outperforming Average Author by 64%. The strong performance of the Writing Summary follows from its reliance on the Author Writing Summary as both a reference for evaluation and a generation constraint. Writing Sheet, despite not being explicitly conditioned on the Author Writing Summary, generalizes well, likely due to its format being the same as the Author Writing Summary. Oracle outperforms Delta and RAG, benefiting from direct conditioning on ground-truth story rules, which enhances adherence to the author’s writing history. However, Oracle underperforms the Writing Sheet and Summary, likely due to its lack of explicit conditioning on the Author Writing Summary. Another possible cause is that the author’s writing style possibly changes over time, which is not fully captured by the ground-truth story rules. See Appendix E.1 for an example.
Reddit shows the highest win-rates among sources:
Reddit exhibits the highest win-rates across all methods, suggesting its stories are more conducive to personalization. This trend likely stems from Reddit’s crowd-sourced writing prompts, where authors create prompts for others. Therefore, there are a range of diverse topics, ranging from power, survival, and mystery to humor and the supernatural. The broad thematic variety may have enabled personalization methods to capture distinct stylistic elements more effectively, resulting in higher win-rates (See the paragraph Themes in Appendix A).
Persona descriptions improve faithfulness to writing history:
We see that adding persona descriptions in the system prompt for Writing Sheet and Summary improves win-rates by around 5% over their nP variants. This result likely follows from personas enhancing the LLM’s role-playing capabilities, allowing for a more faithful representation of the author’s writing style (Jiang et al., 2024; Wang et al., 2024; Wallace et al., 2024).
6.1.2 Similarity to Author Story
@@
Table 3 shows results for similarity to the author story using GPT-4o as LLMstory for our personalization methods. We observe the following:
Writing Sheet and Summary outperform baselines of RAG and Delta:
Oracle achieves the highest win-rates across all methods as its story rules come from the ground-truth author story. Reddit shows the highest Oracle win-rate (91%), followed by 65% on average for other methods, indicating easier personalization likely due to the broad thematic variety in the writing prompts. Writing Sheet and Summary outperform RAG and Delta by 11% and 6%, highlighting the benefit of summarizing the author’s writing history (Richardson et al., 2023).
Persona descriptions help in sources with broad thematic variety in writing prompts:
Writing Summary with persona descriptions outperforms its nP variant by 9% across all sources. However, Writing Sheet shows no significant difference from its nP variant except for Reddit, where it surpasses it by 7%. This result suggests that persona descriptions enhance personalization in sources that are easier to personalize, such as Reddit, but have limited impact elsewhere (Zheng et al., 2024). The broad thematic variety in Reddit’s writing prompts creates a greater divergence between author styles and the Average Story, leading to more informative Author Writing Sheets and persona descriptions that further aid personalization. See Appendix E.2 for an example.
Creativity and Language Use are the easiest categories to personalize:
Figure 2: Win-rate of personalization methods vs. Average Author for Reddit across four narrative categories. D: Delta, S: Writing Summary, WS: Writing Sheet.
Table 10 (in the Appendix) shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024). Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
To better understand personalization performance in Reddit, the easiest source to personalize, Figure 2 shows that Writing Sheet is the best-performing method across all categories, with a 35% advantage over Delta and 20% over Writing Summary in Creativity. Writing Sheet explicitly summarizes differences between the ground-truth author story and the Average Story, providing more informative guidance for personalization. This advantage is particularly evident in Creativity, which is less constrained by the writing prompt. See Appendix E.3 for plots of other data sources.
6.1.3 Traditional Metrics
Table 11 (in the Appendix) shows results with the traditional metrics.
Lexical Overlap and Diversity Metrics yield similar values:
We observe that Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023).
Writing Sheet outperforms other methods, especially for Style metrics:
@@ 6.1.3 Traditional Metrics
We observe that Writing Sheet nP consistently performs best across all metrics, particularly in Style metrics for both Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021). This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization.
6.2 Human Evaluation
Table 3 shows win-rates from human evaluation (Section 5.3) under the Human Eval column, showing that personalization methods—Delta, Writing Sheet, and Summary—outperform the Average Author method by 18% points absolute win-rate in terms of similarity to ground-truth writings. See Appendix G for details on experiment design, and qualitative examples.
Low inter-annotator agreement due to task subjectivity:
Fleiss’ Kappa for inter-annotator agreement ranges from 0.2 to 0.4, indicating low but above-random agreement, likely due to the challenges of annotating long-form subjective texts (Subbiah et al., 2024a). Specifically, different annotators prioritize different narrative categories when determining which story in a pair better reflects the author’s style, leading to disagreement.
Sources like Reddit and New Yorker have better win-rates than others:
Reddit and New Yorker achieve the highest win-rates, with Reddit benefiting from a broad thematic variety in writing prompts and New Yorker from advanced narrative devices like subtext, which humans assess more reliably than LLMs (Subbiah et al., 2024b). In contrast, AO3 and Storium exhibit lower win-rates due to LLMs’ familiarity with fanfiction tropes (AO3) [^3] https://archiveofourown.org/admin_posts/25888 and the lack of strong stylistic differentiation in Storium beyond its open-ended storytelling structure (Xu et al., 2024; Tian et al., 2024). These findings suggest that personalization is most effective when authors engage with diverse topics or possess distinct stylistic traits (such as complex narrative structures, creative elements, and advanced language use). These characteristics, reflected in their past writing history, provide clear signals for us to mimic their writing style.
Narrative categories like Creativity, and Language use are more conducive to personalization:
@@ 6.2 Human Evaluation
Annotators favor personalized stories for Creativity, and Language Use, noting their stronger use of deeper symbolism, thematic richness, layered narratives, and expressive language. However, Average Author was preferred when personalized methods incorrectly introduce elements not present in the author’s ground-truth story, suggesting deviations from the author’s usual style.
@@ 7.1 Personalization and Role-Playing
Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email, abstract, and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024) and 850 for news writing (Shashidhar et al., 2024)). These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025). In contrast, we tackle the more subjective problem of long-form story writing, with author stories averaging 1500 tokens. Unlike prior role-playing approaches that use predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.

### v2

@@ 5 Experiments
We now detail our experimental/evaluation settings.
@@ 5.1 Implementation Details
We prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022) with temperature 0 and a maximum token limit of 4096 for all the methods described above (Section 4) to generate story rules and persona descriptions. Only for story generation (LLMstory) across all methods, including Average Author, we evaluate three different models: GPT-4o [^2] For Reddit, we only consider 5 stories per author for generation totaling 300 stories, due to API cost constraints. (Bubeck et al., 2023), Llama 3.1 8B, and Llama 3.1 70B (Dubey et al., 2024), using a temperature of 0.7 and top_p of 0.95 (Wang et al., 2024) (See Appendix F for results using Llama models). Additionally, we implement an ablation variant of our proposed method for Sheet and Summ that excludes the persona description from the system prompt called Sheet-nP and Summ-nP, respectively.
5.2 Automated Evaluation
We evaluate the performance of personalized story generation using two automated evaluation methods: First, we use LLM-as-a-judge, which is a scalable alternative to human evaluation for complex open-ended text generation tasks (Zheng et al., 2023) and calculate the win-rate of a story generation method against the non-personalized baseline, Average Author, on two aspects:
  • Faithfulness to Writing History: We evaluate how well the generated story aligns with the author history by measuring win-rates using Claims for each narrative category from the Author Writing Summary as the reference (Wang et al., 2023; Yunusov et al., 2024).
@@ 5.2 Automated Evaluation
For both aspects, we prompt OpenAI o4-mini [^3] https://openai.com/index/introducing-o3-and-o4-mini/ in a chain-of-thought manner (Wei et al., 2022) to assign scores (1–5) to both stories. We randomly shuffle the order of stories in comparison, to avoid biases. Scores are broken down into four narrative categories: Plot, Creativity, Development, and Language Use (Saha et al., 2024). For each category, the story with the higher score is declared the winner, and the overall winner is based on the highest total score across all categories; two stories having the same total scores result in a tie. See Appendix D.1 for the prompts. Additionally, we evaluate personalization using traditional metrics (Xie et al., 2023), including lexical overlap, story diversity, and stylistic similarity. See Appendix D.2 for details.
@@ 6 Results and Discussion
We now present experimental results. Our automatic evaluation shows that the proposed methods are effective, particularly for certain sources and narrative categories.
6.1 Faithfulness to Writing History
@@
Table 2 shows results for faithfulness to writing history among GPT-4o-personalized stories. We observe the following:
Method Reddit AO3 Storium N.Mag N.York All
RAG 40 52 42 21 40 39
Delta 56 60 50 64 53 57
Sheet 74 76 73 85 80 78
Sheet-nP 74 80 60 85 86 78
Summ 84 96 87 85 80 87
Summ-nP 76 86 85 100 100 90
Table 2: Win-rates (%) for Faithfulness to Writing History vs. Average Author baseline for GPT-4o. ‘Sheet’ (underlined) performs close to the upper bound established by the ‘Summ’ method and significantly outperforms other personalization methods.
Sheet and Summ achieve best faithfulness to writing history:
The high performance of Summ sets a performance upper bound, since it relies on the Author Writing Summary, the reference for evaluation, to generate the story. Sheet, despite not being explicitly conditioned on the Author Writing Summary, generalizes well (only a 9% gap in win rate compared to Summ), which shows that it is capable in faithfully reflecting the author’s writing history. RAG performs the worst since it relies on a single in-context example to mimic the author’s style. Delta significantly outperforms RAG (by 18%) but significantly underperforms Sheet (by 21%). This result suggests that explicitly summarizing authors’ stylistic characteristics systematically in the form of a Sheet is better than relying on rules extracted from all past author stories. See Appendix E.1.1 and Table 10 for an example.
Persona descriptions benefit amateur sources more than expert sources for Summ:
We see that personas have little impact on Sheet, since the overall win-rates for Sheet and Sheet-nP are identical. This result is likely attributed to using the Author Writing Summary, rather than the Sheet, as the reference for faithfulness evaluation. In contrast, for Summ, adding persona descriptions improves win-rates over Summ-nP for amateur sources such as Reddit, AO3, and Storium, but leads to lower performance for expert sources like the N.Mag and N.York. This result is likely due to personas for expert authors failing to capture the nuanced and subtle characteristics of their writing styles, which are reflected in their Author Writing Summary. See Appendix E.1.2 and Tables 11, 12 for examples.
6.2 Similarity to Author Story
Method Reddit AO3 Storium N.Mag N.York All
Oracle 86 84 78 79 87 83
RAG 38 56 50 43 27 43
Delta 41 60 48 43 33 45
Sheet 62 66 40 71 54 59
Sheet-nP 47 50 33 64 33 45
Summ 45 50 38 50 40 45
Summ-nP 43 53 45 64 20 45
Table 3: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for GPT-4o. Best methods per source are bolded.
@@ 6.2 Similarity to Author Story
Table 3 shows results on similarity to the author story among GPT-4o-personalized stories. We observe the following:
Sheet outperforms all other methods:
The Sheet method achieves the highest overall win-rate of 59%, outperforming Summ by 14%, Delta by 14%, and RAG by 16%. The advantage is more pronounced on Reddit, by far the largest subset of our dataset, where Sheet outperforms Summ by 17%, Delta by 21%, and RAG by 24%. These results suggest that relying on a single in-context example (RAG) is insufficient for capturing personalized writing styles, particularly when authors have more varied profiles, as in Reddit. In contrast, for sources like Storium, where authors have fewer stories and less consistent styles, RAG is more competitive. The Delta method, which uses only story rules, struggles to capture subtle, repeated writing patterns that are better modeled through the longer and more detailed Author Writing Sheet. Finally, Summ’s approach of identifying writing characteristics from all profiling stories at once, without contrasting against Average Author stories, is limited by LLMs’ long-context reasoning capabilities (Shashidhar et al., 2024). See Appendix E.2.1 and Table 13 for an example.
Sheet still lags behind the upper-bound Oracle method:
Although the Sheet method performs best, there remains a substantial overall gap (of 24%) compared to the Oracle method, which highlights the challenging nature of personalization in story generation. This gap likely arises because the Sheet cannot capture all of an author’s writing characteristics within the profiling set, due to generation length constraints (4096 tokens). Moreover, the Sheet does not explicitly rank characteristics based on their relevance to a new writing prompt, which limits its ability to adaptively emphasize the most salient aspects during generation. See Appendix E.2.2 and Table 14 for an example.
Persona descriptions benefit Sheet:
We see that Sheet outperforms its ‘nP’ variant overall by 14%, with a particular gain of 15% on Reddit, the largest subset of our data. Persona descriptions, provided in the system prompt, offer additional control for personalized story generation by summarizing the author’s general writing style independently of the specific writing prompt, complementing the story rules used in the user prompt. See Appendix E.2.3 and Table 15 for an example. In contrast, persona descriptions have limited impact on the Summ method, likely because the Author Writing Summary is less effective than the Author Writing Sheet at capturing subtle author characteristics.
Creativity and Language Use are easier to personalize:
Method Plot Creativity Development Language Use
Oracle 74 82 69 88
Delta 44 55 43 57
Sheet 49 74 51 62
Summ 44 51 50 51
Table 4: Win-rates (%) across narrative categories for Similarity to Author Story, averaged across sources. Best numbers for each method are bolded.
Figure 2: Win-rates proportions across narrative categories for Similarity to Author Story for Reddit. O: Oracle, D: Delta, Sh: Sheet, Su: Summ.
Table 4 shows that all methods have higher respective win-rates on Creativity and Language Use than Plot and Development. This result likely arises because Creativity (e.g., unique elements introduced by the author) and Language Use (e.g., specific diction or use of dialogue) are less dependent on the writing prompt and thus more generalizable across stories (Huot et al., 2024). In contrast, Plot (e.g., conflict introduction or resolution) and Development (e.g., character growth) are closely tied to the writing prompt, making generalization difficult, particularly when there is limited thematic overlap (Tian et al., 2024).
Figure 2 shows that, for our largest story source, Reddit, the Sheet method achieves the highest win-rates across all categories except Development, where it is slightly outperformed by Summ. For Creativity, Sheet performs nearly as well as the Oracle, with only a 5% win-rate gap between them. However, substantial gaps remain between Sheet and Oracle for other categories, which highlights the difficulty of the task and room for future improvement. See Appendix E.2.4 and Figure 22 for plots from other data sources.
6.3 Traditional Metrics
Table 16 in the Appendix shows results with traditional metrics. Overall, we see that the lexical overlap and diversity scores are similar across methods, since all generations use the same LLM, resulting in overlapping lexical distributions that limit these metrics’ sensitivity to stylistic differences (Zheng et al., 2023; Xie et al., 2023).
@@ 6.3 Traditional Metrics
We also see that the Sheet achieves the best stylistic similarity to both Author History and Author Story, measured by LUAR (Rivera-Soto et al., 2021). This result validates the Sheet’s ability to explicitly summarize stylistic deviations from an Average Author, leading to better personalization.
7 Human Evaluation
7.1 Setting
We conduct a human evaluation on a subset of our data to assess narrative subtext and identify insights potentially missed by LLM judges (Chakrabarty et al., 2024; Subbiah et al., 2024b). Annotators compare stories generated by personalization methods (Delta, Sheet, and Summ) against the Average Author method on similarity to the ground-truth author story (Lyu et al., 2024). For each comparison, annotators choose an overall winner by evaluating both stories on all four narrative categories, and provide justifications as free-form responses. Three annotators, recruited via Upwork, evaluated 45 author stories in total. There are 15 shared stories across annotators to assess agreement, and 10 were unique to each annotator to expand coverage (Song et al., 2024). Each author story appeared in three different pairs (one per personalization method), resulting in 135 annotated story pairs. See Appendix G for details on experiment design and the interface (Figure 23).
7.2 Results
Method Reddit AO3 Storium N.Mag N.York All
Delta 44-44 33-33 33-33 33-33 67-11 42-31
Sheet 67-33 44-22 22-33 67-33 78-11 56-27
Summ 44-33 33-44 33-22 44-22 67-22 44-29
Overall 52-37 37-33 30-30 48-30 70-15 47-29
Table 5: Win-rates (%) for Similarity between GPT-4o-generated and Author stories, evaluated by Humans. Each cell (‘X-Y’) shows the method’s (‘X’) and Average Author’s (‘Y’) win-rates, with ties as remainder. Best methods per source are bolded.
Table 5 shows results from our human evaluation. Sheet is clearly the best method, outperforming Average Author by 29% across all sources. On Reddit, our largest source, it reaches a 67% win-rate, further confirming its advantage.
Fair inter-rater and model-human agreement:
Fleiss’ Kappa among human annotators and Cohen’s Kappa between human annotations and the LLM judge (o4-mini) both range from 0.2 to 0.4, indicating fair agreement (Landis and Koch, 1977). This result is likely due to the inherent subjectivity of evaluating long-form texts (Subbiah et al., 2024a), where annotators prioritize different narrative aspects when choosing the overall winner. The scale of human evaluation, constrained by annotation costs (\$639), also contributes to variability. See Appendix G.3 for detailed examples.
Creativity, and Language use are easier to personalize:
@@ 7.2 Results
Same as LLM-as-a-judge, human annotators also favor personalized stories for Creativity and Language Use, noting their stronger use of deeper symbolism, thematic richness, layered narratives, and expressive language. However, Average Author was preferred when personalized methods incorrectly introduce elements not present in the author’s ground-truth story, suggesting deviations from the author’s usual style. See Table 22 in the Appendix for detailed analysis of annotator free-form comments in each narrative category.
Reddit and New Yorker are easier to personalize than other sources:
Reddit and New Yorker show the highest overall win-rates for the personalized methods (52% and 70%, respectively). Reddit benefits from the broad thematic variety of its crowd-sourced writing prompts (see Figure 3 in the Appendix), which leads to diverse author writing styles and makes authors easier to distinguish from the Average Author (see Table 23). New Yorker, with stories written by experts, benefits from the use of advanced narrative devices such as subtext, which humans can evaluate more reliably than LLMs (Subbiah et al., 2024b) (see Table 24). AO3 and Storium show lower overall win-rates for personalized methods. For AO3, LLMs’ familiarity with common fanfiction tropes [^4] https://archiveofourown.org/admin_posts/25888 strengthens the Average Author baseline (see Table 25). For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2024; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 26). These findings suggest that personalization is most effective when authors explore diverse topics or display distinct traits, such as complex narratives, creativity, or advanced language use. Such traits, reflected in past writings, provide clearer signals for personalizing their style.
@@ 8.1 Personalization and Role-Playing
Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024), 850 for news writing (Shashidhar et al., 2024)). These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025). In contrast, we address the more subjective task of long-form story writing, with author stories averaging 1500 tokens. Unlike prior role-playing approaches that rely on predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.